consider the system:
dx/dt = x^2+y
dy/dt = x^2*y^2
Show that, for the solution (x(t),y(t)) with initial ocndition (x(0),y(0)) = (0,1), there is a time t* such that x(t)--> infinity as t--> t*. In other words the solutions blows up in finite time.
1
Expert's answer
2019-10-22T12:13:07-0400

"y(0)=1\ and\ dy/dt = x^2*y^2>=0\implies y(t)\ge 1",

"\implies" "dx/dt \ge x^2+1,\implies"

"\frac{dx/dt}{x^2+1}\ge 1 \implies"

"(arctan(x(t)))'_t>=1,\implies"

"arctan(x(t))\ge x(0)+t=0+t=t,"

Therefore there are bounded increasing sequence of points

"{t_n},\ (t_n<\pi/2), t_n\to t^*\le\pi/2:"

"arctan(x(t_n))\to \pi/2\implies"

"x(t_n)\to\infty",when "t_n\to t^*."